S. Lee Enterprises produces two models of lawn tractor: SL1 and SL2. The models are both produced in the company’s Louisville factory and go through identical assembly operations. The difference is in the quality (hence, cost) of the parts used. The following data are available for November, the latest month for which information is available:

SL1 SL2 Total
Number assembled 1,450 1,950 3,400
Materials cost per tractor $ 1,900 $ 2,400
Other costs:
Direct labor $ 2,200,000
Supplies 680,000
Other overhead 3,580,000


Required:

S. Lee Enterprises uses operations costing and assigns conversion costs based on the number of units assembled. Compute the cost of each model assembled in November.

SL1 SL2 Total $

Operation Cost

Materials Cost

Total Cost

Number of Units

Unit Cost
Answer #1
SL1 SL2 Total
Operation Cost 2755000 3705000 6460000
Materials Cost 2755000 4680000 7435000
Total Cost 5510000 8385000 13895000
Number of Units 1450 1950 3400
Unit Cost 3800 4300 8100
Workings:
Total conversion cost 6460000 =2200000+680000+3580000
Divide by Total units assembled 3400
Conversion cost per unit 1900
Conversion cost allocated:
SL1 2755000 =1450*1900
SL2 3705000 =1950*1900
Materials costs:
SL1 2755000 =1450*1900
SL2 4680000 =1950*2400